THE PURE ROTATIONAL SPECTRA OF VN (X ^{3}Δ_{r}) and VO (X ^{4}Σ^{−})

\underline{M. A. FLORY} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectra of VN (X ^{3}Δ_{r}) and VO (X ^{4}Σ^{−}) in their ground electronic states have been recorded in the frequency range 295-525 GHz. In each case, seven rotational transitions were measured. Fine structure and vanadium hyperfine splittings were resolved in these data. These radicals were produced by reacting gas-phase VCl_{4} with N_{2} (VN) or residual O_{2} (VO) in an AC discharge. Rotational, fine, and hyperfine constants were determined. In particular, the hyperfine constants a, b, and c for VN have been established from a global fit to the spectra. The parameters obtained are in good agreement with previous studies.